OpenStudy (anonymous):

If sinθ = 8/17 and tan θ> 0 find the exact value of sin2θ

OpenStudy (anonymous):

Help please

OpenStudy (dls):

Sin theta=8/17 Sin2 theta=2 sin theta cos theta Also,sin^2 theta+ cos^2 theta=1 From here cos theta= sqrt(1- sin^2 theta) So find cos theta and put everything in the formula :)